OpenStudy (anonymous):

Dave eats 2/5 of a box cereal every day. In how many days will he finish 4 boxes of cereal?

OpenStudy (anonymous):

4=4/1=20/5 (20/5)/(2/5)= your answer

OpenStudy (anonymous):

where you get 20 from?

OpenStudy (anonymous):

\[4=\left[\begin{matrix}4 & *5 \\ 1 & *5\end{matrix}\right]=\left(\begin{matrix}20 \\ 5\end{matrix}\right)\]

OpenStudy (anonymous):

In 1 day he eats 2/5 and he needs to eat 4 boxes or 10 times that x/5 = 4 x equals 2 times how many days it takes him. 20/5 = 4 20 / 2 = 10 days.